# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

### v2

@@ 3.3 Trails of the signal in CCF maps
TOI 1518 b is a misaligned planet, as presented in Cabot et al. (2021). A nodal precession analysis made by Watanabe et al. (2024) underline a change of the impact parameter of the planet in function of time (from b = 0.91497 in 2019 to b= 0.8797 in 2022). However, this change does not affect significantly our result and the geometry of the system in our short timescale so we decided to keep the value of Cabot et al. (2021) (b=0,9036) as reference for our study. The geometry of the system is presented in Fig. 16. Fig. 5 presents the iron cross-correlation map on the stellar rest frame on the left and shifted to the planetary rest frame on the right. The dashed yellow line on the right underlines the expected position of the planet if the atmosphere is static. As the planet is misaligned, the Doppler shadow effect (already observed in Cont et al. 2021) only affects the planetary signal at the beginning of the transit and we have decided to not mask it or remove it. This is a different case from WASP-76 b (Pelletier et al., 2023; Ehrenreich et al., 2020) where it has to be removed as the Doppler shadow crosses the planetary trail.
The iron track shown in Fig. 5 is slightly shifted from the theoretical planetary velocity computed with the orbital parameters of Table 2. Previous high-resolution observations have demonstrated the ability to resolve time variations of this atmospheric track (Ehrenreich et al., 2020; Borsa et al., 2021). We thus investigated this with our data. We binned the CCF similarly to Wardenier et al. (2024) to increase the planetary signal. We divided the 25 in-transit frames observed in both datasets into nine bins. We use scipy.optimize.curve fit to fit a Gaussian to each bin between 10 km/s (in the planetary rest frame). This gave better results than performing the fit across a broader range of velocities. The results are shown in Fig. 6. The velocity center of the atmospheric track changes with time (Fig. 6, Left Panel), becoming more blue-shifted from around +1 km.s-1 to around -8 km.s-1. This is similar to what happens in the case of other UHJs like WASP-76b and WASP-121b (Ehrenreich et al., 2020; Borsa et al., 2021). The signal amplitude varies, with the function of the orbital phase being more important at mid-transit than at the beginning or end of the transit.
We further detect the time-varying trace of the CCF for six different species: Fe, Fe+, Ca, Ca+, Na, Mg (Fig. 7). All species show a qualitatively similar behavior to the Fe trace, with a blueshift over time.
@@ 3.4 Discussion
The detection of VO was made with a significance of 4.9 sigma using the newly released HyVO line list this year (Bowesman et al., 2024). VO is notoriously difficult to detect in exoplanet atmospheres; however, Pelletier et al. 2023 demonstrates the feasibility of such detections when employing a more accurate line list. For WASP-76b, the VOmyt line list was useful for the detection of VO. Fig. 8 illustrates that this line list was unsuccessful in retrieving VO signals in the observations of TOI-1518b. This discrepancy can be attributed to a stronger VO signal in WASP-76b, along with the use of three transit observations compared to only two for TOI-1518b. In this study, we demonstrate the superior capability of the HyVO line list, which successfully detected VO (see Fig. 8). The detected signal is blueshifted by as much as -12.11 0.22 km/s (see Table 3), which is significantly higher than the maximum blueshift observed for other species in TOI-1518b, recorded at -5.16 0.57 km/s. Currently, it remains uncertain whether this difference is due to a shift in the line list itself or a physical shift in the atmosphere, potentially stemming from the varying localization of VO compared to metals and ions in the atmosphere of TOI-1518b. The presence of strong optical absorbers such as VO and TiO in the atmosphere of UHJs is crucial for thermal inversion phenomena (Fortney et al., 2008). This finding underscores the significance of utilizing more accurate line lists for the detection of these molecules, especially in cases where the signals may be weaker than those observed in WASP-76b. Further studies employing this new line list could reveal the presence of VO in targets where it was previously undetectable, as highlighted in Borsa et al. (2021).
Detecting Fe/Fe+, Ca/Ca+, Ti/Ti+, and potentially V/VO raises questions. Due to their presence in specific parts of the atmosphere, these species should not co-exist. For example, Fe+ is expected more in the hotter dayside atmosphere, whereas Fe should be more present on the cooler limbs and nightside. However, the tracks of Fe and Fe+ seen in Figure 7 show a similar blueshifting trend, meaning that they are likely probing similar atmospheric regions, albeit probably at different pressures.
@@
Several maps, such as Fe, Mg, Ca, or Cr, show a parasitic signal at km/s and km/s. This signal is a residual of the Rossiter-McLaughlin effect observed in the CCF map in Fig. 5. The Rossiter-McLaughlin effect presents an anti-correlation signal at negative . Therefore, it is not shown here, but the residual positive correlation visible in yellow around the anti-correlation signal in the CCF map is the parasite signal observed in the diagram mentioned previously. Possible biases due to Rossiter Mc-Laughing effect are present at phase below -5 degree and are represented in grey on the trail map of Fig. 7. The positive positive correlation at phases above 5 degree is far from the planetary signal 5 ( km/s). For the case of Mg, a signal is also visible at expected but = 100 km.s-1. This is due to a strong Fe+ line in the Mg triplet. This is also visible in the CCF maps of both species, where the residual signal of the other can be observed as highlighted in Fig. 18.
Fig. 20 highlights some limitations of the Gaussian profiles used to parameterize the iron trail of TOI-1518 b as most one-dimensional CCF do not follow a simple Gaussian profile. Then we decided to center the fitted Gaussian profile on the maximum of the 1D-CCF even if this resulted in a misestimation of the FWHM and probably of the error bars of the measured V. Another uncertainty might be due to the PCA applied to the data. Fig. 19 presents the iron trail of TOI-1518 b with different numbers of principal components removed. The square root of the variance of the Doppler shifts across all numbers of removed components, PCA has been added to the uncertainty quoted in the covariance matrix of the Gaussian fit obtained from scipy.optimize.curve fit. The PCA does not change the transit trend even when many components are removed for all species except Ca+ (Fig.21). The PCA is essential for detecting faint species. To maintain consistency, we removed the same number of components (three) from all species, even for those where it might not have been necessary. The only exception is Ca+, as the PCA significantly impacts the signal, even when fewer components are removed, due to its very high signal strength. Therefore, we decided not to apply PCA to Ca+ in order to achieve a more robust analysis of its trail. To understand the physics behind these six time-resolved absorption, a comparison with Global Circulation Models (GCMs) is conducted in Sect. 4.
@@ 4 Comparison with Global Circulation Models
The 3D nature of UHJs can explain the trace of iron in the CCF map observed in this work. We compare the iron signal’s blueshift with two different 3D global circulation models (GCMs) scenarios to interpret the data.
@@ 4.1 Model description
We consider five SPARC/MITgcm models of TOI 1518 b, three are presented in Fig. 9. The SPARC/MITgcm was initially introduced by Showman et al. (2009).
It has been widely used to study the atmospheric physics and chemistry of (ultra-)hot Jupiters (Fortney et al., 2010; Showman et al., 2013; Kataria et al., 2013; Parmentier et al., 2018; Tan et al., 2024). Our models of TOI-1518 b are based on work by Tan et al. (2024). We explore the effect of atmospheric drag (Showman et al., 2013; Komacek & Showman, 2016; Parmentier & Crossfield, 2018), in considering four models with drag timescales from s (strong drag) s (weak drag). The drag timescale encompasses a wide range of physical mechanisms, including turbulent mixing (Li & Goodman, 2010), Lorentz-force braking of ionized winds within the planet’s magnetic field (Perna et al., 2010a), and Ohmic dissipation (Perna et al., 2010b).
These GCMs account for heat transport due to H2 dissociation and recombination (e.g., Bell & Cowan (2018); Komacek & Tan (2018);Tan & Komacek (2019); Roth et al. (2021)). H2 thermally dissociates on the dayside, after which atomic hydrogen gets advected to the nightside, where it recombines into H2 and releases latent heat. When the atmospheric circulation is predominantly eastward, most of this heat is dumped on the evening limb, resulting in a temperature asymmetry between the eastern and western regions of the atmosphere (first and second column in Fig.9).
Drag restores energy to the atmosphere. Increasing drag strength (i.e lowering ) slows down winds in the atmosphere hindering this model’s heat transport. In strong drag cases, the temperature structure becomes symmetric, resulting in similar chemical compositions in the morning and evening limbs. (last column in Fig. 9). In the strong-drag model, there is only a day-to-night flow as the equatorial jet gets suppressed. Table 4 summarizes some other important parameters of the two SPARC/MITgcm models. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer. All models were run at a horizontal resolution of C32, corresponding to roughly 128 cells in longitude and 64 in latitude. Before computing phase-dependent spectra of the GCMs with gCMCRT, we bin the outputs down to 32 latitudes and 64 longitudes, as in Wardenier et al. (2021, 2023, 2024).
@@
| Orbital Period | 1,6442 s (1.903 days) |
| Pressure range | 200 2 |
| Radius at bottom | 1.3405 108m (1,875 R) |
@@
| Radiative transfer | non-grey (see ) |
@@ 4.2 Injection and Cross-Correlation maps
We compute phase-dependent transmission spectra of the two GCM models across the MAROON-X spectral range (between 490 and 920 nm) using gCMCRT (Lee et al., 2022). The calculations account for Doppler shifts due to planet rotation and winds. Section 2 of Wardenier et al. (2023) shows the radiative transfer and post-processing details. Therefore, we will only briefly summarize the gCMCRT setup for TOI-1518 b analysis. Before feeding the GCM outputs into gCMCRT, we map the atmospheric structures onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. We account for the fact that each atmospheric column has a different scale height set by local gravity, temperature, and mean molecular weight. For each of the five TOI-1518 b models, we simulate 25 spectra (equidistant in orbital phase) between phase angles 8 degrees, covering the in-transit part of our observations. We assume an orbital impact parameter of 0.9036, a semi-major axis of 0.039 au, a stellar radius of 1.95 R⊙, and an orbital period of 1.90 days. At each orbital phase angle, gCMCRT simulates a transmission spectrum by randomly shooting photon packets at the part of the planet limb that is blocking the star and evaluating the optical depth encountered by each photon packet. Due to the geometry of TOI-1518 b (Fig. 16), the illumination differs from the case of WASP-121 b where the impact parameter is near 0 (Fig. 10, Right Panel). In this calculation, the code accounts for Doppler shifts imparted on the opacities by the radial component of the local wind vector and planet rotation (Wardenier et al., 2021). The transit depth at a specific wavelength is calculated by averaging it over all photon packets. To accurately represent the shapes, depths, and shifts of the spectral lines, we use 105 photon packets per wavelength. Since we do not explicitly account for scattering, the direction of propagation for the photon packets remains constant throughout the calculation. The spectra are calculated at a resolution of , which differs from Wardenier et al. (2023, 2024). We include the same set of continuum opacities and line species for the radiative transfer as in Wardenier et al. (2023). To properly compare the GCM spectra with the data, we need to inject the GCMs into the data and perform PCA, similar to what we did for the data. The different steps were as follows:
Inject the spectra at each orbital phase of the observations at the expected Doppler shift calculated with the expected of Equation 1.
@@
Do the cross-correlation with the same Fe template and methods used for Sect. 3.1.
@@
Figs. 11 and 12 show the CCF signals of Fe and Fe+ that we obtain for each of the models, with the real data plotted on top. We interpolated the cross-correlation results over the same phase grid as for the observations of the two transits. We then performed the same analysis using the fit with a Gaussian profile, which allowed us to extract the position, amplitude, and FWHM of the model’s signal.
@@ 4.3 Discussion
Fig. 10 shows the importance of considering the impact parameter in calculating the spectra. We clearly see a big difference in the spectrum for the same GCM output of the SPARC/MITgcm. The blueshift is way more pronounced in the case of a high-impact parameter. The jump in the signal driven by planet rotation also occurs earlier in the case of a high-impact parameter (around phase = -2.5 deg) than in the case of impact parameter b =0, where it happens at mid-transit. This is because the illuminated part of the planet differs in the two cases. When b=0, the leading limb is illuminated at the start of the transit, both limbs are illuminated in the middle, and the trailing limb is illuminated at the end. In contrast, only the southern parts of the limbs are illuminated in the case of b =0.9.
Fig.11 (Left Panel) shows the Fe trails of the five GCMs simulations, along with the Doppler shifts of Fe measured in section 3.1. The absorption trails stand between the stronger drag models ( and s). The weak drag and drag-free models produce Doppler shifts that are too much blueshifted due to strong winds that are not attenuated by drag effects.
Fig.12 now compares the GCM outputs to the Fe+ trace. There is a pretty good match between the observation and the stronger drag model (s in purple), especially since the GCMs are not optimized to fit the data. Overall, the fact that both Fe and Fe+ trails are very similar points towards a strong drag scenario, where the iron is fully condensed on the nightside and both Fe and Fe+ probe similar regions close to the limb.
@@
Fig.13 shows the trails for the 6 species of Fig.7 with the signals of the stronger drag models. For Ca, Na, Mg, and Ca+ the signal straddles between both these GCMs, as in the case of Fe. The strong drag scenario is then preferred also by the other species detected
However the distribution of ionized iron (Fe+), which is present on the entire day side, and neutral iron (Fe), found only at the limbs and in the inner part (Fig.9), may account for the variations in the models and data shape of the trail. In particular, the drag-free and weak drag models exhibit a Fe signal influenced primarily by strong wind effects, with velocities ranging from around -5 to -12 km/s. The wind in the two stronger drag models is almost nullified at the Fe localization. There is a clear difference between the ingress and egress parts of the transit, which is caused by the redshift and blueshift of the planet’s rotation (Fig.11). The wind at the location of Fe+ in the upper atmosphere is less affected in the strong drag model, leading to a more continuous trail between ingress and egress part of the transit (Fig.12).
This effect is more extreme in the case of the Ca+ signal, which is redshifted over compared to the models and the other trails. Ionised species appear then more redshifted than neutral species. The models used in this study assume consistent drag in the atmosphere; however, this finding highlights the potential variation of drag with altitude. We expect neutral species to dominate in the lower part of the atmosphere, while ionized species will be more prevalent in the upper layers. This variation in drag could be attributed to an increase in magnetic drag with altitude, as demonstrated in previous studies (e.g., Beltz et al. 2022).
The agreement with an intense drag model atmosphere was already observed in the case of WASP-121 b by Wardenier et al. 2024 by comparing with ESPRESSO data at with similar models. Moreover, as in WASP-76 b (e.g., Wardenier et al. 2021; Savel et al. 2022), the Fe trail of TOI-1518 b provides clear evidence of a strong thermochemical asymmetry between the morning and evening terminator. Fig. 11, we can see that in the first half of the transit, the data are more blueshifted than the models with stronger drags (between phases at -5 and 0 degrees), which probably implies that the signal from the trailing limb is stronger (compared to a model with no asymmetry)
Then, a similar scenario to the one observed in WASP-76 b (Fig. 2 in Wardenier et al. 2021) seems to appear where, at the beginning of the transit, the Fe signal is mainly influenced by the leading limb, where the redshift from the planet’s rotation balances out the blueshift from day-to-night winds. The Doppler shift should become more negative as the trailing limb comes into view because the planet’s rotation and day-to-night winds cause a blueshift in the signal.
Looking at Fig. 11, the Doppler-shift measurement of iron (Fe) in the third bin of the transit (just before mid-transit) may be considered an outlier, but this could also be attributed to a limitation of the Gaussian profile used in this context. In Fig. 10, the CCF of the GCM exhibits a double-peak structure with one peak near 0 km/s and another near -7 km/s. The data signal does not seem to reflect this double-peak scenario, instead showing a non-Gaussian signal, as discussed in Sect. 3.4 (see Fig. 20).
The amplitude of the signal shown in the middle panel of Fig.11 and Fig.12 is mainly influenced by the portion of the planet that transits in front of the star, depicted by the yellow dashed line. All models follow this trend even though it is more robust in the second part of the transit compared to the data. The signal amplitude does not underline any asymmetry between the two limbs or between the models because of the specific geometry of TOI-1518 b. It still interesting to note that the amplitude of the signal in the Fe simulations is lowering with stronger drag effect due to the double peak scenarios underlined before. This double peak scenario appear because of wind reduced by the drag as discussed above. For the Fe+ simulations, the winds is higher and so the signal is smoother with all the models following the same trend due to the fraction of the planet.
The FWHM is also in agreement with all models in the first order. This parameter is not constrained by the approach of a simple Gaussian fit as presented in Fig.20, and another approach would be more efficient, but this is not the primary purpose of this paper. The fact that the models have FWHM less important with drag effect increasing may be due to the double-peak structure present from phase -5 degrees to +2.5 degrees, corresponding to the lower FWHM observed. As our simple Gaussian profile only fits one of the two peaks, there may be an underestimation of the FWHM of the model signal compared to the data signal, which does not emphasise this double peak shape.

### v3

@@ 3.4 Trails of the signal in CCF maps
Fig. 5 presents the time-variation of the cross-correlation function calculated with an iron template and shifted into the planetary rest frame. The dashed yellow line on the right underlines the expected position of the planet if the atmosphere is static and homogeneous. As the planet is misaligned, the Doppler shadow effect (already observed in Cont et al. 2021) only affects the planetary signal at the beginning of the transit and we have thus decided to not mask it nor to remove it.
The iron track shown in Fig. 5 is slightly shifted from the theoretical planetary velocity computed with the orbital parameters of Table 2. Previous high-resolution observations have demonstrated the ability to resolve time variations of this atmospheric track (Ehrenreich et al., 2020; Borsa et al., 2021). We thus investigated this with our data. We binned the CCF similarly to Wardenier et al. (2024) to increase the planetary signal. We divided the 25 in-transit frames observed in both datasets into nine bins. We use scipy.optimize.curve_fit to fit a Gaussian to each bin between 10 km/s in the planetary rest frame. This gave better results than performing the fit across a broader range of velocities. The results are shown in Fig. 6. The velocity center of the atmospheric track changes with time (Fig. 6, Left Panel), becoming more blue-shifted from around +1 km.s-1 to around -8 km.s-1. This is similar to what happens in the case of other UHJs like WASP-76b and WASP-121b (Ehrenreich et al., 2020; Borsa et al., 2021). The signal’s amplitude reaches its peak during mid-transit, rather than at the beginning or the end.
We further detect the time-varying trace of the CCF for six different species: Fe, , Ca, , Na, Mg (Fig. 7). All species show a qualitatively similar behavior to the Fe trace, with a blueshift over time.
@@ 3.5 Discussion
The detection of VO was made with a significance of 4.9 sigma using the newly released HyVO line list this year (Bowesman et al., 2024). VO is notoriously difficult to detect in exoplanet atmospheres; however, Pelletier et al. 2023 demonstrates the feasibility of such detections when employing a more accurate line list. For WASP-76b, the VOmyt line list was useful for the detection of . Fig. 8 illustrates that this line list was unsuccessful in retrieving VO signals in the observations of TOI-1518 b. This discrepancy can be attributed to a stronger signal in WASP-76b, along with the use of three transit observations compared to only two for TOI-1518 b. In this study, we demonstrate the superior capability of the HyVO line list, which successfully detected VO (see Fig. 8). The detected signal is blueshifted by as much as -12.11 0.22 km/s (see Table 3), which is significantly higher than the maximum blueshift observed for other species in TOI-1518 b, recorded at -5.16 0.57 km/s. Currently, it remains uncertain whether this difference is due to a shift in the line list itself or a physical shift in the atmosphere, potentially stemming from the varying localization of VO compared to metals and ions in the atmosphere of TOI-1518 b. The presence of strong optical absorbers such as VO and TiO in the atmosphere of UHJs is crucial for thermal inversion phenomena (Fortney et al., 2008). This finding underscores the significance of utilizing more accurate line lists for the detection of these molecules, especially in cases where the signals may be weaker than those observed in WASP-76b. Further studies employing this new line list could reveal the presence of VO in targets where it was previously undetectable, as highlighted in Borsa et al. (2021).
Detecting , , , and potentially raises questions. For example, is expected more in the hotter dayside atmosphere, whereas should be more present on the cooler limbs and nightside. However, the tracks of and seen in Figure 7 show a similar blueshifting trend, meaning that they are likely probing similar atmospheric regions.
@@
Several maps, such as Fe, Mg, Ca, or Cr, show a parasitic signal at km/s and km/s. This signal is a residual of the Rossiter-McLaughlin effect observed in the CCF map in Fig. 5. The Rossiter-McLaughlin effect presents an anti-correlation signal at negative . Therefore, it is not shown here, but the residual positive correlation visible in yellow around the anti-correlation signal in the CCF map is the parasite signal observed in the diagram mentioned previously. Possible biases due to Rossiter Mc-Laughing effect are present at phase below -5 degree and are represented in grey on the trail map of Fig. 7. The positive positive correlation at phases above 5 degree is far from the planetary signal 5 ( km/s). For the case of Mg, a signal is also visible at expected but = 100 km.s-1. This is due to a strong line in the Mg triplet. This is also visible in the CCF maps of both species, where the residual signal of the other can be observed as highlighted in Fig. 20.
Fig. 22 highlights some limitations of the Gaussian profiles used to parameterize the iron trail of TOI-1518 b as most one-dimensional CCF do not follow a simple Gaussian profile. Then we decided to center the fitted Gaussian profile on the maximum of the 1D-CCF even if this resulted in a misestimation of the FWHM and probably of the error bars of the measured V. Another uncertainty might be due to the PCA applied to the data. Fig. 21 presents the iron trail of TOI-1518 b with different numbers of principal components removed. The square root of the variance of the Doppler shifts across all numbers of removed components, PCA has been added to the uncertainty quoted in the covariance matrix of the Gaussian fit obtained from scipy.optimize.curve_fit. The PCA does not change the transit trend even when many components are removed for all species except (Fig.23). The PCA is essential for detecting faint species. To maintain consistency, we removed the same number of components (three) from all species, even for those where it might not have been necessary. The only exception is , as the PCA significantly impacts the signal, even when fewer components are removed, due to its very high signal strength. Therefore, we decided not to apply PCA to in order to achieve a more robust analysis of its trail.
@@ 4 Comparison with global circulation models
To understand the physics behind the six time-resolved absorption trails we observed, we now compare our data with a suite of global circulation models.
@@ 4.1 Model description
We consider five SPARC/MITgcm models of TOI 1518 b, three are presented in Fig. 9. The SPARC/MITgcm was initially introduced by Showman et al. (2009). It has been widely used to study the atmospheric physics and chemistry of (ultra-)hot Jupiters (Fortney et al., 2010; Showman et al., 2013; Kataria et al., 2013; Parmentier et al., 2018). Our cloud-free models of TOI-1518 b are based on work by Tan et al. (2024). As in previous works (Showman et al., 2013; Komacek & Showman, 2016; Parmentier & Crossfield, 2018), we parameterize several source of dissipation (including turbulent mixing(Li & Goodman, 2010), shocks (Heng, 2012) , hydrodynamic instabilities (Fromang et al., 2016), or magnetic drag(Perna et al., 2010; Beltz et al., 2022)) through a Newtonian relaxation term in velocity applied to the momentum equation. The relaxation timescale, called the drag timescale hereafter, ranges from s (strong drag) s (weak drag).
These GCMs account for heat transport due to H2 dissociation and recombination (e.g., Bell & Cowan (2018); Komacek & Tan (2018);Tan & Komacek (2019); Roth et al. (2021)). H2 thermally dissociates on the dayside, after which atomic hydrogen gets advected to the nightside, where it recombines into H2 and releases latent heat. When the atmospheric circulation is predominantly eastward, most of this heat is pumped on the evening limb, resulting in a temperature asymmetry between the eastern and western regions of the atmosphere (first and second column in Fig.9). Drag restores energy to the atmosphere. Increasing drag strength (i.e lowering ) slows down winds in the atmosphere hindering this model’s heat transport. In strong drag cases, the temperature structure becomes symmetric, resulting in similar chemical compositions in the morning and evening limbs. (last column in Fig. 9). In the strong-drag model, there is only a day-to-night flow as the equatorial jet gets suppressed (Showman et al., 2013). Table 4 summarizes some other important parameters of the two SPARC/MITgcm models. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer (which include species such as and driving thermal inversion on the planet dayside). All models were run at a horizontal resolution of C32, corresponding to roughly 128 cells in longitude and 64 in latitude. Before computing phase-dependent spectra of the GCMs with gCMCRT, we bin the outputs down to 32 latitudes and 64 longitudes, as in Wardenier et al. (2021, 2023, 2024).
@@
| Orbital Period | 1.6442 s (1.903 days) |
| Pressure range | 200 2 bars |
| Radius at bottom | 1.3405 108m (1.875 R) |
@@
| Radiative transfer | non-grey (see Kataria et al. 2013) |
@@ 4.2 Injection and Cross-Correlation maps
We compute phase-dependent transmission spectra of the five GCM models across the MAROON-X spectral range (between 490 and 920 nm) using gCMCRT (Lee et al., 2022). The calculations account for Doppler shifts due to planet rotation and winds. Section 2 of Wardenier et al. (2023) shows the radiative transfer and post-processing details. Therefore, we will only briefly summarize the gCMCRT setup for TOI-1518 b analysis. Before feeding the GCM outputs into gCMCRT, we map the atmospheric structures onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. We account for the fact that each atmospheric column has a different scale height set by local gravity, temperature, and mean molecular weight. For each of the five TOI-1518 b models, we simulate 25 spectra (equidistant in orbital phase) between phase angles 8 degrees, covering the in-transit part of our observations. At each orbital phase angle, gCMCRT simulates a transmission spectrum by randomly shooting photon packets at the part of the planet limb that is blocking the star and evaluating the optical depth encountered by each photon packet. Due to the geometry of TOI-1518 b (Fig. 18), the part of the planet that is illuminated differs from the case of WASP-121 b where the impact parameter is near 0 (Fig. 10, Right Panel). In this calculation, the code accounts for Doppler shifts imparted on the opacities by the radial component of the local wind vector and planet rotation (Wardenier et al., 2021). The transit depth at a specific wavelength is calculated by averaging it over all photon packets. To accurately represent the shapes, depths, and shifts of the spectral lines, we use photon packets per wavelength. Since we do not explicitly account for scattering, the direction of propagation for the photon packets remains constant throughout the calculation. The spectra are simulated at a native resolution of R = 300 000, and then convolved down to the instrument resolution of , which differs from Wardenier et al. (2023, 2024). We include the same set of continuum opacities and line species for the radiative transfer as in Wardenier et al. (2023). To properly compare the GCM spectra with the data, we need to inject the GCMs into the data and perform PCA, similar to what we did for the data. The different steps were as follows:
Inject the spectra at each orbital phase into the three components that were removed during the data reduction process, using the velocity value from Equation 1.
@@
Do the cross-correlation with the same Fe template and methods used for Sect. 3.2.
@@
We then interpolated the cross-correlation results over the same phase grid as for the observations of the two transits. We then performed the same analysis using the fit with a Gaussian profile, which allowed us to extract the position, amplitude, and FWHM of the model’s signal.
Figs. 11 and 12 show the CCF signals of and that we obtain for each of the models, with the real data plotted on top.
@@ 4.3 Discussion
Fig.11 (Left Panel) shows the Fe trails of the five GCMs simulations, along with the Doppler shifts of measured in section 3.2. All models show a signal that is blueshifting with time. This behaviour was previously observed by Ehrenreich et al. (2020); Borsa et al. (2021) and discussed in many papers (Wardenier et al., 2021, 2023; Savel et al., 2022; Beltz et al., 2022, 2023). Because the planet rotates during the transit, the start of the transit probes the dayside of the leading limb and the nightside of the trailing limb. It is thus dominated by the hotter leading limb. The planet rotation red-shifts the leading limb by 3.5 km/s, which is compensated by the blue shift of the day-to-night winds. In the second part of the transit, the opposite happens and the signal is dominated by the trailing limb, which is blueshifted by rotation. The change from being dominated by the leading and the trailing limb during transit is responsible for the blueshifting of the signal with time.
In the case of TOI-1518 b’s atmosphere the high impact parameter reinforces this blue shift. Indeed, as shown in Fig. 10, because the planet is close to being grazing, we almost never probe the full atmospheric limb. The signal from the first half of the transit is therefore entirely due to the leading limb, with only a very small contribution to the trailing limb, and the opposite is true for the second part of the transit. As a consequence, we find that a given planet is expected to have a stronger blueshift with time when observed with a high impact parameter than if it has a low impact parameter.
Furthermore, the global shift of the planetary trace is linked to the day-to-night winds that blueshift both leading and trailing hemisphere signals. These day-to-night winds are strongly affected by drag. As shown in Fig.11, models without drag produce an iron signal that is too blueshifted compared to the observations. Overall, the planetary signal stands between the stronger drag models ( and s). When looking at the amplitude and at the FWHM of the signal, however, we see that the s model is favoured, because the s provides a signal that is too small and not wide enough compared to data.
If we now look at the trace (Fig.12), we see that the s overestimates the blue shift, whereas the s ( in purple) is a much better match. For the amplitude of the model is unaffected by the drag and the FHWM errorbars are too large to differentiate between the two strong drag models. The apparent contradiction between the Fe trace favouring the s model and the trace favouring the stronger, s drag model could be an indication of the presence of ohmic drag in the atmosphere. Indeed, as shown by Beltz et al. (2022), Ohmic drag strength strongly depends on planetary location, being stronger in the hotter dayside and at lower pressures. Because the equilibrium is temperature dependent, the signal naturally probes more into the hotter planetary dayside than the Fe signal, where the temperatures are hotter and magnetic drag effects are expected to be larger.
@@
Fig.13 shows the trails for the 6 species of Fig.7 with the signals of the stronger drag models. For , , , and the signal straddles between both these GCMs, as in the case of Fe. We therefore confirm that the strong drag scenario is preferred by the signal of all the other species. A similar agreement with an intense drag model atmosphere was already observed in the case of WASP-121 b by Wardenier et al. 2024, pointing out that strong drag is common in ultra-hot Jupiter atmospheres.
The trail of the is significantly different from the ones of the other species, as it starts strongly red-shifted. This is clearly not captured by the model. We calculated the contribution functions of our model, and find that the CCF of most species (including Ca, Fe and ) probe the and bar layers whereas the probes layers up to . We fitted directly the resolved individual lines using a method similar to that of Prinoth et al. (2024) and found a transit depth corresponding to an effective radius of . This value is similar to the planetary Roche radius (R at Lagrange Point 1 following Gu et al. (2003)). As a consequence, we believe that the lines likely probe the outflow of the planet. This region shows a different atmospheric flow than the deeper regions, with an important contribution from planet-star interactions, not modeled in our GCM. Our results for are similar to the recent observations of H-alpha for WASP-121b (Seidel et al., 2025).
The CCF of the GCM, presented in Fig. 10, exhibits a double-peak structure with one centered near 0 km/s and another around -7 km/s. This feature arise from the contribution of both atmospheric limb. Fitting this complex signal with a simple Gaussian profile is inherently limiting, as a single-peaked function cannot accurately capture the asymmetric nature of the CCF. Parameters such as the FWHM and amplitude may be misrepresented, potentially leading to an incomplete interpretation of the atmospheric dynamics. Finally, for vanadium oxide, the predictions from the GCMs do not match the observed signal observed in Fig. 8. The GCMs estimate a change in Kp of about 5 km/s and a Vres) ranging from -6 to -13 km/s (decreasing with decreasing drag) when the signal observed in the data is at around -33 km/s and Vres around -13km/s.